Run Summary: Arizona Partners, January 5 - 10, 2011

The mask exchange mechanism of Lucifer-1 failed just before the start of the AZ run, and the decision was made to cancel spectroscopic observations, which unfortunately made up the bulk of the observations we hoped to do. A call for proposals for back-up imaging with Lucifer and LBC was made, and several projects were sent it, on very short notice.

Weather was poor for most of the run.

LUCIFER-specific

LBC-specific suggestions

1. The LBC does not function reliably. It should be fixed, or retired. During this run it failed in a major way at least once every 24 hours: the blue filter wheel malfunctioned, requiring about 8 hours of repair; the housekeeping pressure and temperature alarms stopped the camera from functioning for no reason; the Red ccd controller failed in the middle of the 2nd or 3rd OB of the first night we tried observing. Data was lost, telescope time was lost. Cycling power can take as long as 2 hours. LBC's level of reliability is not acceptable for an instrument on a major facility like the LBT. One should be able to run the instrument all night without a failure, at least once in a while.

I think it would be good to fix LBC, as a backup for times when the AO secondaries fail. Until it's fixed, I'm going to recommend that Arizona not award time to use it, since we have access to other large format imagers at other facilities.

2. Fix RB_science to run correctly on bias images; Fix WCS parameters in bias images (several key parameters are currently "nan" in the fits header)

Bias frames taken in the afternoon are an important check that LBC is working.

We took bias images, and tried running RB_science to display them simultaneously. RB_science should be used by all partners, (as opposed to Red_science and Blue_science) -- we modified the .cshrc for AZ to use it per instructions from Olga; need to describe in the documentaion. The problem is that the bias images were not displaying all 4 chips and there was only a single frame.

Michael found the problem in the RB_science code. It seems that the images are being loaded in 'Mosaic WCS' mode, and that the appropriate header keywords are not in LBC bias frames for these to be loaded appropriately.

There are two possible straightforward fixes:

1) RB_science should look in the image header. If it is a 'bias' frame, or otherwise does not have appropriate WCS information, then the images should be opened in 'Mosaic IRAF' mode rather than 'WCS Mosaic' mode.

2) The real fix would be if ALL LBC images have appropriate header information so that they can be loaded in 'Mosaic IRAF' mode.

3. Make it obvious when the camera has failed and data is not being taken. When the red controller stopped, the OB continued taking blue data, and the only indication was the image count, buried in the gui. Need a flashing red light or something to alert the observer when the camera fails.
Topic revision: r2 - 10 Jan 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback